General Meeting Summary 9/10/20

From Cuawiki
Jump to: navigation, search

PARTICIPANTS: Aaron Brown, Alexandre Camsonne, Arshak Asaturyan, Bogdan Wojsekhowski, Chris Cuevas, Hamlet Mkrtchyan, Jacob Murphy, Steve Wood, Vardan Tadevosyan, Vladimir Berdnikov, Charles Hyde, Marie Boer, Petr Stepanov, Rolf Ent, Steven Lassiter, Hakob Voskanyan,


SUBSYSTEM UPDATES

FRAME (Carlos)

  • Work ongoing - continue assembly and testing of all components


HV CONNECTOR - HV RATING CONCERNS (Chris, Steve)

  • DSG raised the concern that HV ratings seem low for some items
  • Based on FEG initial analysis it is fine to move forward - need to prepare a report, might look in some more detail at specific items, perhaps perform local tests
  • Actions:
  • Near term: Brief and concise report from Electrical to allow to move forward (Chris, Fernando, et al.)
  • Longer term: When begin experiment need safety documentation and procedure for NPS

Notes on OSP:

  • coverage by protective box
  • interlock back to CAEN supply - max operating voltage
  • identify any potential touch points - might implement zero voltage check point (indicator light or similar to indicate that there are exposed parts inside the box)
  • might need to define voltage rating criterion - typical safety ratings use general criteria, not application specific


DSG UPDATES (Aaron)

  • HV cables: 750/1100 fabricated
  • Control and monitoring screens
  • Interlock system - need feedback on expected radiation
  • Temperature monitoring - could check with accelerator and/or target groups on suitable sensors
  • Actions:
  • need feedback on expected radiation
  • check with ACC and/or target groups on sensors
  • RG cable and connector - provide details to Steve Wood


DESIGN/MECHANICAL (Steven Lassiter for Paulo)

  • Reqs being prepared and submitted for structural and hardware for NPS platform
  • Waiting on detector center of mass drawings - for making the lifting frame
  • Plan to check out G0 rails: surface properties, count
  • One final analysis of magnet platform: HB plates modification (for multiple installations/deinstallations)


CRYSTALS/REFLECTOR SHAPING (Vladimir)

  • Crystals from Crytur coming in every month - visual inspection, so far all looks good
  • Reflector shaping tool
  • Josh working on drawings
  • Machine shop open for critical tasks
  • Initial discussion of stacking of crystals and strategy for NPS, TCS/NPS, FCAL, EIC
  • one model is SICCAS on outside and Crytur inside
  • Actions:
  • count number crystals needed for all devices
  • table of where need highest performance crystals
  • radiation dose of all experiments (DVCS/SIDIS: Ho San's studies, also need WACS and TCS)


HV DIVIDER

  • Action: check on status with Fernando


TCS PLANS (Marie)

  • Discussion of the list of items developed to demonstrate that experiment can be successful, e.g. identify particles of interest
  • Further updates and discussions will follow


NEXT MEETING: THURSDAY 1 OCTOBER AT 9AM (ET)